# VGDiffZero: Text-to-image Diffusion Models Can Be Zero-shot Visual Grounders

### 14 Dec 2023 v2

@@ 1 Introduction
Recently, large-scale vision-language pre-trained models [1, 2] have demonstrated strong performance across a wide range of downstream tasks. Among them, most vision-language (VL) tasks can be broadly categorized into two types - generative tasks and discriminative tasks. For VL generative tasks, text-to-image generative diffusion models, such as Stable Diffusion [3], have shown their powerful ability to generate high-fidelity and diverse images from text descriptions. To be specific, these generative diffusion models are firstly pre-trained on large-scale text-image pairs datasets like LAION-5B [4] to learn the correlations between textual descriptions and visual concepts, then applying them in VL generative tasks, such as image editing [5] and inpainting [6]. For VL discriminative tasks, the predominant approaches are to obtain vision-language alignment abilities from large-scale pre-training datasets, which can then be utilized to facilitate downstream visual-language reasoning tasks [7, 8]. Among various VL discriminative tasks, visual grounding [9] is one of the most challenging one, in which the aim is to localize a target object in an image given a textual description. Most supervised visual grounding works [10, 11, 12] study how to effectively fuse cross-modality features extracted independently by each encoder. Though these methods have achieved good performance on general visual grounding benchmarks, such as RefCOCO and RefCOCO+ [9], the training and fine-tuning cost is prohibitively expensive and time-consuming. Moreover, collecting task-specific data is even more challenging since it requires accurate descriptions of the target area as well as high-quality bounding box annotations [13].
@@ 3.1 Experimental Settings
Datasets and Evaluation Metrics. We evaluate VGDiffZero on the RefCOCO [9], RefCOCO+ [9] and RefCOCOg [18], which are common benchmarks for visual grounding, collected from MS-COCO [24]. The RefCOCO, RefCOCO+, and RefCOCOg datasets contain 19,994, 19,992, and 26,771 images with 142,210, 141,564, and 104,560 referring expressions, respectively. RefCOCO and RefCOCO+ have shorter expressions (avg 1.6 nouns, 3.6 words), while RefCOCOg has longer, more complex expressions (avg 2.8 nouns, 8.4 words). Following previous works [10, 22], we use Accuracy@0.5 as the evaluation metric to assess the performance of our proposed method and other baseline models.
Baselines. We compare VGDiffZero with related methods: Random, CPT-Blk [22], and three diffusion-based baseline methods. These are explained in detail as: (1) Random: Randomly selecting a detected proposal as the prediction. (2) CPT-Blk [22]: A strong zero-shot visual grounding baseline that shades detected proposals with different colors and uses a masked language prompt in which the referring expression is followed by “in [MASK] color”. The color with the highest probability by the pre-trained masked language model VinVL [7] is chosen as prediction. (3) Cropping: Isolating detected proposals by cropping the image around each proposal, then passing the cropped images through VGDiffZero. (4) Masking: Similar to Cropping, but using masking to isolate detected proposals before passing through VGDiffZero. (5) VGDiffZero w/ Single IPM: Given two sets of prediction errors from isolating proposal methods (IPM), cropping and masking, selecting the proposal with the lowest prediction error as the final prediction.
@@ 3.2 Experimental Results and Analysis
Quantitative Comparison. The main experimental results on RefCOCO, RefCOCO+, and RefCOCO are reported in  Table 1, from which we observe that: (1) Our proposed VGDiffZero outperforms other zero-shot visual grounding baseline methods. Notably, on the RefCOCO+ test A, VGDiffZero achieves 5.79% higher accuracy compared to CPT-Blk [22]. This indicates that VGDiffZero can effectively and directly leverage the vision-language alignment abilities learned by pre-trained VL generative models to perform well on visual grounding task. (2) Masking to isolate object proposals achieves superior performance compared to cropping on most datasets, which suggests that preserving the global location of proposals plays a role in the visual grounding task. (3) Using both masking and cropping to isolate proposals achieves higher accuracy than using either method alone. This indicates that considering both the global and local contexts of isolated proposals enables more robust performance on visual grounding tasks.
@@ 4 Conclusions
In this work, we propose a simple yet effective zero-shot visual grounding framework VGDiffZero, which directly leverages the pre-trained vision-language alignment abilities of text-to-image diffusion models. Through the designed comprehensive region-scoring method, our VGDiffZero can consider both the global and local contexts of each isolated object proposal. Extensive experimental results demonstrate that VGDiffZero achieves satisfactory performance on three general visual grounding benchmarks.

### 02 Jan 2024 v3

@@ 1 Introduction
Recently, large-scale vision-language pre-trained models [1, 2] have demonstrated strong performance across a wide range of downstream tasks. Among them, most vision-language (VL) tasks can be broadly categorized into two types - generative tasks and discriminative tasks. For VL generative tasks, text-to-image generative diffusion models, such as Stable Diffusion [3], have shown their powerful ability to generate high-fidelity and diverse images from text descriptions. To be specific, these generative diffusion models are firstly pre-trained on large-scale text-image pairs datasets like LAION-5B [4] to learn the correlations between textual descriptions and visual concepts, then applying them in VL generative tasks, such as image editing [5] and inpainting [6]. For VL discriminative tasks, the predominant approaches are to obtain vision-language alignment abilities from large-scale pre-training datasets, which can then be utilized to facilitate downstream visual-language reasoning tasks [7, 8, 9]. Among various VL discriminative tasks, visual grounding [10] is one of the most challenging one, in which the aim is to localize a target object in an image given a textual description. Most supervised visual grounding works [11, 12, 13] study how to effectively fuse cross-modality features extracted independently by each encoder. Though these methods have achieved good performance on general visual grounding benchmarks, such as RefCOCO and RefCOCO+ [10], the training and fine-tuning cost is prohibitively expensive and time-consuming. Moreover, collecting task-specific data is even more challenging since it requires accurate descriptions of the target area as well as high-quality bounding box annotations [14].
@@ 3.1 Experimental Settings
Datasets and Evaluation Metrics. We evaluate VGDiffZero on three widely-used VG benchmarks: RefCOCO [10], RefCOCO+ [10] and RefCOCOg [19]. RefCOCO, RefCOCO+, and RefCOCOg datasets contain 19,994, 19,992, and 26,771 images with 142,210, 141,564, and 104,560 referring expressions, respectively. RefCOCO and RefCOCO+ have shorter expressions (avg 1.6 nouns, 3.6 words), while RefCOCOg has longer, more complex expressions (avg 2.8 nouns, 8.4 words). We use Accuracy@0.5 as the evaluation metrics.
Baselines. We compare VGDiffZero with related methods: (1) Random: Randomly selecting a detected proposal as the prediction. (2) CPT-Blk [23]: A strong zero-shot visual grounding baseline that shades detected proposals with different colors and uses a masked language prompt in which the referring expression is followed by “in [MASK] color”. The color with the highest probability by the pre-trained masked language model VinVL [7] is chosen as prediction. (3) Cropping: Isolating detected proposals by cropping the image around each proposal, then passing the cropped images through VGDiffZero. (4) Masking: Similar to Cropping, but using masking to isolate detected proposals before passing through VGDiffZero. (5) VGDiffZero w/ Single IPM: Given two sets of prediction errors from isolating proposal methods (IPM), cropping and masking, selecting the proposal with the lowest prediction error as the final prediction.
@@ 3.2 Experimental Results and Analysis
Quantitative Comparison. The main experimental results are reported in Table 1, from which we can observe that: (1) Our proposed VGDiffZero outperforms other zero-shot visual grounding baseline methods. Notably, on the RefCOCO+ test A, VGDiffZero achieves 5.79% higher accuracy compared to CPT-Blk [23]. This indicates that VGDiffZero can effectively and directly leverage the vision-language alignment abilities learned by pre-trained VL generative models to perform well on visual grounding task. (2) Masking to isolate object proposals achieves superior performance compared to cropping on most datasets, which suggests that preserving the global location of proposals plays a role in the visual grounding task. (3) Using both masking and cropping to isolate proposals achieves higher accuracy than using either method alone. This indicates that considering both the global and local contexts of isolated proposals enables more robust performance on visual grounding tasks.
@@ 4 Conclusions
In this paper, we propose VGDiffZero, a novel zero-shot visual grounding framework that leverages pre-trained text-to-image diffusion models’ vision-language alignment abilities. Through the designed comprehensive region-scoring method, our VGDiffZero can consider both the global and local contexts of each isolated object proposal. Extensive experimental results demonstrate that VGDiffZero achieves satisfactory performance on three general visual grounding benchmarks.
Acknowledgements. This research was supported by STI 2030—Major Projects (2022ZD0208800), NSFC General Program (Grant No. 62176215), and NSFC Young Scientists Fund (Grant No. 62001316).